# Consistency-based Abductive Reasoning over Perceptual Errors of Multiple Pre-trained Models in Novel Environments

### v4

@@ Introduction
Technical Preliminaries. We consider the problem of object identification for a given set of perception data $\Omega$ (Cheng and Han 2016) and assume the availability of a set $\eta$ of perception models $\mathcal{F}=\{f_{1},f_{2},\dots,f_{\eta}\}$, where each model generates a set of predictions over a shared set of $m$ object classes $\mathcal{C}=\{c_{1},...,c_{m}\}$. Under the unique name assumption [^1] See Section 1 of the appendix for the implementation details of how we employ this assumption., every object $\omega\in\Omega$ identified by one or more models has at least one associated fact of the form $f_{i}(\omega)=c_{j}$, meaning that model $i$ has identified object $\omega$ as belonging to class $c_{j}$. We will denote the set of these facts (hereafter “observations”) as $O$.
@@
Intuitively, if metacognitive cue is present for object $\omega$ and it was classified as class $c$ by model $f_{i}$, then we can assume that there was an error made in the assignment. In (Kricheli et al. 2024), the metacognitive cues (referred to as “conditions” in that work) are selected from a set of candidate conditions. We include details of how we applied their learning algorithm in the appendix.
@@
As observable in Figure 4, while there is a general trend of decreasing F1-scores for all methods as the average environmental intensity increases—indicating the inherent difficulty of operating in more severe novel environments—our IP+TB approach (represented by red diamonds) consistently achieves the highest F1-scores across the entire spectrum of intensities. In nearly all instances, IP+TB surpasses the Best Individual Model, and significantly outperforms the Average Models, Majority Vote, and our Heuristic Search (HS+TB) approach in terms of F1-score. This consistent superiority, irrespective of the environmental intensity level, underscores the effectiveness of IP+TB in robustly integrating model predictions and managing inconsistencies. For instance, even at higher intensity levels where all methods experience performance degradation, IP+TB maintains a clear advantage. The same analysis for Accuracy shows comparable trends, and is provided in the appendix.
@@
As expected, increasing $\epsilon$ generally leads to higher precision but lower recall, demonstrating the inherent trade-off controlled by this parameter. Notably, this stricter filtering also tends to reduce the baseline level of inconsistency among the remaining predictions. This analysis highlights how tuning $\epsilon$ shapes the pool of candidate predictions subsequently processed by our IP+TB and HS+TB abduction methods aiming to maximize valid assignments under the global inconsistency constraint $\delta$. Detailed sensitivity plots for all test sets are provided in the appendix.
Hyperparameter Sensitivity. We analyzed the sensitivity of our approaches to the maximum inconsistency threshold hyperparameter $\delta$, which applies to IP+TB and HS+TB, and the internal recall parameter $\epsilon$ used within the IP+TB optimization. Figure 5 (right) illustrates this analysis for the MDS-A_1 test set. The 3D plot of IP+TB show that the F1 score and accuracy reach their maximum value in the ranges of $\delta$ between 0.1 and 0.3, and for $\epsilon$ in 0.1 and 0.5. Similarly, the performance of HS+TB (center) stabilizes quickly as $\delta$ increases above a small initial value (e.g., 0.2), which is associated with the maximum level of inconsistency present in the initial test set (Figure 5, left); detailed sensitivity plots for all test sets are provided in the appendix.
Running Time Analysis and Complexity. A key practical difference between our approaches is computational cost. Our empirical results, illustrated in a running time plot available in the appendix, align with the theoretical complexities of the methods. As expected, our polynomial-time HS+TB approach is significantly faster than the exact IP+TB method. While the IP solver’s running time is higher due to its exact nature, it remained tractable for the instances explored here. We observed that the average processing time per object for IP+TB did not increase steeply with more objects. We hypothesize that this is due to a combination of amortized solver overheads and structural dataset characteristics that allow for faster resolution. A detailed study of these issues is future work.

### v6

@@ Introduction
Technical Preliminaries. We consider the problem of object identification for a given set of perception data $\Omega$ (Cheng and Han 2016) and assume the availability of a set $\eta$ of perception models $\mathcal{F}=\{f_{1},f_{2},\dots,f_{\eta}\}$, where each model generates a set of predictions over a shared set of $m$ object classes $\mathcal{C}=\{c_{1},...,c_{m}\}$. Under the unique name assumption [^1] See Section 1 of the appendix in the extended version for the implementation details of how we employ this assumption., every object $\omega\in\Omega$ identified by one or more models has at least one associated fact of the form $f_{i}(\omega)=c_{j}$, meaning that model $i$ has identified object $\omega$ as belonging to class $c_{j}$. We will denote the set of these facts (hereafter “observations”) as $O$.
@@
Intuitively, if metacognitive cue is present for object $\omega$ and it was classified as class $c$ by model $f_{i}$, then we can assume that there was an error made in the assignment. In (Kricheli et al. 2024), the metacognitive cues (referred to as “conditions” in that work) are selected from a set of candidate conditions. We include details of how we applied their learning algorithm in extended version.
@@
As observable in Figure 4, while there is a general trend of decreasing F1-scores for all methods as the average environmental intensity increases—indicating the inherent difficulty of operating in more severe novel environments—our IP+TB approach (represented by red diamonds) consistently achieves the highest F1-scores across the entire spectrum of intensities. In nearly all instances, IP+TB surpasses the Best Individual Model, and significantly outperforms the Average Models, Majority Vote, and our Heuristic Search (HS+TB) approach in terms of F1-score. This consistent superiority, irrespective of the environmental intensity level, underscores the effectiveness of IP+TB in robustly integrating model predictions and managing inconsistencies. For instance, even at higher intensity levels where all methods experience performance degradation, IP+TB maintains a clear advantage. The same analysis for Accuracy shows comparable trends, and is provided in the extended version.
@@
As expected, increasing $\epsilon$ generally leads to higher precision but lower recall, demonstrating the inherent trade-off controlled by this parameter. Notably, this stricter filtering also tends to reduce the baseline level of inconsistency among the remaining predictions. This analysis highlights how tuning $\epsilon$ shapes the pool of candidate predictions subsequently processed by our IP+TB and HS+TB abduction methods aiming to maximize valid assignments under the global inconsistency constraint $\delta$. Detailed sensitivity plots for all test sets are provided in extended version.
Hyperparameter Sensitivity. We analyzed the sensitivity of our approaches to the maximum inconsistency threshold hyperparameter $\delta$, which applies to IP+TB and HS+TB, and the internal recall parameter $\epsilon$ used within the IP+TB optimization. Figure 5 (right) illustrates this analysis for the MDS-A_1 test set. The 3D plot of IP+TB show that the F1 score and accuracy reach their maximum value in the ranges of $\delta$ between 0.1 and 0.3, and for $\epsilon$ in 0.1 and 0.5. Similarly, the performance of HS+TB (center) stabilizes quickly as $\delta$ increases above a small initial value (e.g., 0.2), which is associated with the maximum level of inconsistency present in the initial test set (Figure 5, left); detailed sensitivity plots for all test sets are provided in the extended version.
Running Time Analysis and Complexity. A key practical difference between our approaches is computational cost. Our empirical results, illustrated in a running time plot available in the extended version, align with the theoretical complexities of the methods. As expected, our polynomial-time HS+TB approach is significantly faster than the exact IP+TB method. While the IP solver’s running time is higher due to its exact nature, it remained tractable for the instances explored here. We observed that the average processing time per object for IP+TB did not increase steeply with more objects. We hypothesize that this is due to a combination of amortized solver overheads and structural dataset characteristics that allow for faster resolution. A detailed study of these issues is future work.